OpenStudy (anonymous):

What is the value of the expression? 9(8) ______ 9(-4) * 9(10)

OpenStudy (anonymous):

@Leong @stacy.reed08 @dan815

OpenStudy (anonymous):

you mean 9 times 8 over [9 times (-4)] times [ 9 times 10] ?

OpenStudy (sh3lsh):

\[\frac{ 9*4*2 }{9*4*2*-1*9*5}\]

OpenStudy (sh3lsh):

Can you solve it from there?

OpenStudy (anonymous):

yes, one sec....

OpenStudy (sh3lsh):

Hint. Like terms cancel out. So it's just 1/-45

OpenStudy (anonymous):

yeah, that is what i got

OpenStudy (sh3lsh):

bam you're the bomb

OpenStudy (mathstudent55):

\(\dfrac{9(8)}{9(-4) \times 9(10)} =\) \(=\dfrac{72}{-36 \times 90} \) \(=\dfrac{\cancel{72}~2}{\cancel{-36}~-1 \times 90} \) \(= -\dfrac{2}{90} \) \(= -\dfrac{1}{45} \)

OpenStudy (anonymous):

@sh3lsh : is 9 times (-4), not -8 so you can't : 9*2*4

OpenStudy (mathstudent55):

@Leong @sh3lsh did it correctly. The 2 in 9 * 2 * 4 in the denominator comes from the 10 which is 2 * 5. This is how he did it: \(\dfrac{9(\color{red}{8})}{9(\color{blue}{-4}) \times 9(\color{green}{10})} =\) \(=\dfrac{9 \times \color{red}{4 \times 2}}{9 \times \color{blue}{(-1) \times 4} \times 9 \times \color{green}{2 \times 5}} \) \(=\dfrac{9 \times \color{red}{4 \times 2}}{9 \times \color{blue}{4} \times \color{green}{2 \times \color{blue}{(-1) } \times 9 \times 5}} \) \(=\dfrac{\cancel 9 \times \cancel 4 \times \cancel 2 ~1}{\cancel 9 \times \cancel 4 \times \cancel 2 \times (-1) \times 9 \times 5} \) \(= \dfrac{1}{-45} \) \(=-\dfrac{1}{45} \)

OpenStudy (anonymous):

ohhhhhhh